OpenStudy (anonymous):

Express 1/(4-i)^2 in cartesian form.. is the answer 15/289 + 8i/289?? Am i rite??

razor99 (razor99):

@Mandy_Nakamoto what kind of equation is that

OpenStudy (anonymous):

it says to change it to x+iy form.. under the complex number topic..

razor99 (razor99):

i think your are correct lol not sure

OpenStudy (anonymous):

oh you have to draw it on Argand plane I think

OpenStudy (hoblos):

yes you are right XD

OpenStudy (anonymous):

yeah but Argand Plane will only give a graphical solution

razor99 (razor99):

yaaa katrina kaif to the resque

OpenStudy (hoblos):

1/(4-i)^2 = 1/(15 - 8i) multiply by the conjugate ( 15+8i) (15+8i)/ 289 = 15/289 + 8i/289

OpenStudy (anonymous):

Yay! im rite!!!! huhuhu...

razor99 (razor99):

good one Mandy

OpenStudy (anonymous):

But about the argand plane.. i have to change the equation to cartesian form first rite??

OpenStudy (anonymous):

no in order to draw it on argand plane you only need the x + iy format and the point is (x,y) understand

OpenStudy (anonymous):

So bascially the point you will get is (15/289 , 8/289)

razor99 (razor99):

wow Rajarshi you are awesome at maths

OpenStudy (anonymous):

but isnt cartesian form=x+iy form??

OpenStudy (anonymous):

 get it